OpenStudy (anonymous):

The amount of heat needed to raise 2.0 kg of a substance by 80 K is 33 kJ. What is the specific heat of the substance? A. 210 (J/kg*K) B. 2,100 (J/kg*K) C. 1,300 (J/kg*K) D. 1.3 * 10^6 (J/kg*K)

OpenStudy (michele_laino):

it is simple, we have to apply this formula: \[c = \frac{1}{m}\frac{{\Delta Q}}{{\Delta t}} = \frac{1}{2}\frac{{33 \times {{10}^3}}}{{80}}\]

OpenStudy (anonymous):

ok! so we get this? 206.25? so our solution should be choice A?

OpenStudy (michele_laino):

yes! correct!

OpenStudy (anonymous):

yay!! thank you!! :D

OpenStudy (michele_laino):

thanks!! :)